OpenStudy (anonymous):

********medals and fans********At their going-out-of-business sale, Bill's Bikes sold 39 bikes and had 50 left. Which equation could be used to find x, the number of bikes the store had originally?

OpenStudy (anonymous):

x = no. of bikes originally x= 50 +39

OpenStudy (anonymous):

89? not a chose x + 39 = 50 x - 50 = 39 x + 50 = -39 x - 39 = -50

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Its A

OpenStudy (anonymous):

nope I got it wrong